when mike was 10 years old, his father was 34 years. Now his father is twice the age of Mike . What is Mikes age?
Answer by Maths68(1474)

Let
Present age of Mike = m
Present age of Father = f


when mike was 10 years old, his father was 34 years.
or father is 24 years older than Mike
f=m+24......(1)
Now his father is twice the age of Mike
f=2m......(2)
Put the value of f in (1)
2m=m+24
2m-m=24
m=24
Put the value of m in (1)
f=24+24
f=48

Present age of Mike = m = 24 years old
Present age of Father = f = 48 years old